3. In MMW class, a student’s quiz scores are 89, 92, 82, and 77. What score does the student need on the fifth quiz 

to have a mean quiz score of 85? All the quizzes have equal weights. Round your answer to the nearest whole 

number.



1
Expert's answer
2021-11-11T15:08:57-0500

Quiz scove is 89%, 92%, 82% and 77%

We want to find scove of "5^t" quiz so that the mean scove= 85%

Assume that fifth quiz have scove X

"Mean= \frac{Sum \; of \; all \; quiz \; scoves }{Number \; of\; quizes} \\

0.85=\frac{0.89 + 0.92 + 0.82 + 0.77 + X}{5} \\

0.85=\frac{3.4 +X}{5} \\

4.25=3.4 +X \\

X=4.25 -3.4 \\

X=0.85 \\

X=85\; \%"